# Pricing Options on Forwards in Function-Valued Affine Stochastic Volatility Models

### v1

@@ Abstract.
We study the pricing of European-style options written on forward contracts within function-valued infinite-dimensional affine stochastic volatility models. The dynamics of the underlying forward price curves are modeled within the Heath-Jarrow-Morton-Musiela framework as solution to a stochastic partial differential equation modulated by a stochastic volatility process. We analyze two classes of affine stochastic volatility models: (i) a Gaussian model governed by a finite-rank Wishart process, and (ii) a pure-jump affine model extending the Barndorff–-Nielsen–Shephard framework with state-dependent jumps in the covariance component. For both models, we derive conditions for the existence of exponential moments and develop semi-closed Fourier-based pricing formulas for vanilla call and put options written on forward price curves. Our approach allows for tractable pricing in models with infinitely many risk factors, thereby capturing maturity-specific and term structure risk essential in forward markets.
@@ 1. Introduction
We consider a tractable framework for modeling forward price and forward rate curves in commodity and fixed-income markets, and develop semi-closed pricing formulas for the valuation of European vanilla options written on (potentially) the entire forward curve. Following the Heath–Jarrow–Morton–Musiela (HJMM) paradigm, originally proposed for interest-rate markets by [29] and later transferred to commodity markets by [11, 3], we describe forward prices directly in time-to-maturity coordinates. In this representation, the forward curve evolves according to a hyperbolic stochastic partial differential equation (SPDE) driven by a cylindrical Brownian motion whose covariance is itself modeled by a stochastic process that we call the instantaneous covariance process. We investigate two affine specifications of the covariance process: (i) An extension of the finite-rank Wishart process to Hilbert–Schmidt operators introduced in [14] and (ii) a state-dependent jump extension of the operator-valued Ornstein–Uhlenbeck process introduced in [16]. As both models are from the affine class, their associated Fourier-Laplace transforms admit semi-explicit expressions in terms of the solutions to a set of generalized Riccati equations, which allows for semi-closed pricing formulas for vanilla call and put options. The derivation of these pricing formulas, together with rigorous conditions for the existence of exponential moments and a numerical analysis, constitut the principal contribution of this article.
@@ 1.1. The HJMM framework
For any time $t\geq 0$ and maturity date $T\geq t$, the forward price $F(t,T)$ represents the agreed-upon price at time $t$ for a transaction occurring at the future date $T$. Utilizing the Musiela parametrization $x\equiv T-t$, forward prices become functions of time-to-maturity. We define the logarithmic forward prices as $f_{t}(x)\equiv\ln(F(t,t+x))$, where $x,t\geq 0$, and call the mapping $x\mapsto f_{t}(x)$ the (logarithmic) forward price curve at time $t\geq 0$.
The forward curve dynamics are described by the stochastic partial differential equation (SPDE):
$\begin{cases}\mathop{}\!\textnormal{d}f_{t}(x)&=\left(\frac{\partial}{\partial x}f_{t}(x)+g_{t}(x)\right)\mathop{}\!\textnormal{d}t+\sum_{i=1}^{d}\sigma_{t}^{(i)}(x)\mathop{}\!\textnormal{d}W_{t}^{(i)},\\ f_{0}(x)&=F(0,x),\end{cases}$ (1.1)
@@
The functions $\{\sigma_{t}^{(i)}(x))_{0\leq t\leq T}$, ${i=1},\ldots,{d}$, are referred to as the instantaneous volatilities of the forward curve. They characterize the local covariance structure between forward prices of different maturities. The integer $d\in\mathbb{N}\cup\{+\infty\}$ is called the rank of the noise and corresponds to the number of independent risk factors driving the forward curves. If $d<\infty$, we refer to (1.1) as a finite-rank HJM model; otherwise, it is an infinite-rank HJM model.
@@ 1.2. Stochastic covariance models for forward curve dynamics
The diffusion term in (1.1) can be interpreted in a more abstract form as the stochastic integral of an operator-valued process $(\sigma_{t})_{0\leq t\leq T}$ acting on the increments of a cylindrical Brownian motion $(\mathrm{d}W_{t})_{0\leq t\leq T}$. We understand $(\sigma_{t})_{0\leq t\leq T}$ as the stochastic volatility operator, encoding the time-varying covariance structure of the forward curve dynamics. The functions $\sigma_{t}^{(i)}$ can therefore be interpreted as the principal components of $\sigma_{t}$, capturing the volatility structure along different directions in the underlying space of forward curves. The joint evolution of the forward curve and its stochastic volatility, represented by the process $(f_{t},\sigma_{t})_{0\leq t\leq T}$, thus defines an infinite-dimensional stochastic volatility model for forward curve dynamics.
The motivation for adopting a stochastic volatility framework stems from empirical evidence in commodity markets: volatility is time-varying, exhibits clustering, and is prone to abrupt spikes; see, e.g., [24, 6]. Classical models with constant volatility cannot capture these phenomena, whereas stochastic volatility models allow for a more realistic and flexible description of market behavior. Furthermore, to accurately capture volatility smiles in commodity option markets, incorporating stochastic volatility is essential. In contrast to approaches that perform a priori dimension reduction of the covariance, we do not restrict the rank of the volatility operator based on principal component analysis (PCA) or other variance-based criteria. Instead, we adopt a flexible, full-rank modeling approach, retaining all potential sources of risk. This is motivated by the observation that discarding components solely on the basis of explained variance can result in suboptimal choices for higher-order, non-linear functionals, such as those relevant for pricing, hedging, or risk management [12]. Note that even when the true rank is only moderately high, our methodology first emphasizes maximal expressiveness and generality. Finite-rank approximations are introduced a posteriori, and we study convergence of option prices as the rank increases. Empirical studies support this design choice: for instance, [1] and [34] demonstrate that power forward markets exhibit substantial idiosyncratic risk, and that high-dimensional noise is needed to capture the observed variability. In particular, the latter shows that even in PCA, around 10 factors are typically required to explain the majority of the variance, underlining the importance of accommodating high-dimensional dynamics from the outset. Here our function-valued formulation helps, as it is essentially non-parametric, relying on the full initial forward curve and the full covariance operator as model inputs. The covariance structure is encoded via a positive kernel operator, which can capture empirical features with a minimal set of hyperparameters and enables direct calibration to market data, including option prices and historical term structure dynamics.
@@ 1.3. Contribution
In this paper, we contribute to this growing literature by developing a general pricing framework for European options written on forward contracts under function-valued affine stochastic volatility. We formulate the forward curve dynamics within the HJMM framework and model the instantaneous covariance as an operator-valued affine process. Two concrete model classes are considered: a Gaussian specification driven by an infinite-dimensional Wishart process, and a pure-jump model based on an extension of an operator-valued Ornstein–Uhlenbeck process to accommodate state-dependent jumps. The inclusion of jumps in the volatility process enhances the model’s ability to reflect sudden market shocks, which commonly arise from supply disruptions, geopolitical tensions, or abrupt changes in demand; see, e.g. [35]. In both settings, we rigorously extend the affine transform formula to a complex strip in an infinite-dimensional Hilbert space and establish conditions under which exponential moments exist in real and complex domains. In particular, we characterize the maximal domain of the moment generating function using generalized Riccati equations in Banach spaces, building on the analytic machinery introduced in [36], and extending prior finite-dimensional results such as those in [33].
Based on these results, we derive semi-closed-form pricing formulas for vanilla call and put options, which enable efficient numerical evaluation via Fourier inversion. We validate the accuracy and robustness of this pricing methodology through a comparison with Monte Carlo simulation. The numerical results show that both the moment-generating function (MGF) and option prices derived via the affine transform are consistent with those obtained from Monte Carlo simulations, see Figure 1, 2 and 3 below. This comparison not only validates the accuracy of our approach but also illustrates its computational advantages. For example, in the case of the infinite-dimensional Wishart-type volatility model, direct simulation is challenging due to the absence of a closed-form expression for the square root of the operator-valued volatility process. Approximating this square root requires advanced numerical techniques, making Monte Carlo simulation computationally intensive. Similarly, in the case of the infinite-dimensional pure-jump affine volatility model, Monte Carlo methods demand significant computational resources, both in memory and runtime, due to the complexity of the jump structure. By contrast, the affine structure allows for efficient computation of key quantities such as the MGF and derivative prices without resorting to discretization of the infinite-dimensional dynamics. This leads to considerable gains in computational speed and accuracy. These advantages make affine stochastic volatility models particularly suitable for practical applications. In particular, they facilitate the calibration to implied volatility surfaces and option price curves observed in fixed-income or commodity markets, which is a notoriously delicate task. Our results demonstrate that, despite the infinite-dimensional nature of the models, their tractability and analytical structure make them powerful tools for capturing the rich dynamics in forward and futures markets.
@@ 1.4. Layout of the article
The remainder of the paper is organized as follows. In Section 2, we introduce the modeling framework for forward curve dynamics using the HJMM approach and define the class of infinite-dimensional stochastic volatility models under consideration. Sections 2.1 and 2.2 present two concrete affine specifications for the volatility process: a pure-jump model based on an extension of an operator-valued Ornstein–Uhlenbeck process, and a Gaussian model driven by an infinite-dimensional Wishart process. In Section 3, we provide semi-closed-form pricing formulas for European call and put options on forwards, obtained via Fourier inversion. Section 4 presents a numerical study comparing the affine pricing formulas with Monte Carlo simulation methods. We highlight the accuracy, robustness, and computational efficiency of the affine approach, especially in settings where simulation becomes numerically challenging. Finally, in Section 5, we study the existence of the exponential moments.
@@ 1.5. Notation
For $(X,\tau)$ a topological space and $S\subset X$ we let $\mathcal{B}(S)$ denote the Borel-$\sigma$-algebra generated by the relative topology on $S$. We denote by $C^{k}([0,T];S)$ the space of $S$-valued $k$-times continuously differentiable functions on $[0,T]$. Throughout this article we fix a separable, infinite-dimensional real Hilbert space $(H,\langle\cdot,\cdot\rangle_{H})$. The complexification of a real Hilbert space is the complex Hilbert space $H^{\mathbb{C}}=H\oplus_{\mathbb{R}}\mathrm{i}H=\{x+\mathrm{i}y\colon x,y\in H\}$ equipped with the complex inner product $\langle x+\mathrm{i}y,u+\mathrm{i}v\rangle_{H^{\mathbb{C}}}=\langle x,u\rangle_{H}-\langle y,v\rangle_{H}+\mathrm{i}\langle y,u\rangle_{H}+\mathrm{i}\langle x,v\rangle_{H}$, for all $x,y,u,v\in H$. For $z=x+\mathrm{i}y\in H^{\mathbb{C}}$, we say that $x=\Re(z)$ and $y=\mathcal{I}(z)$ are the real and imaginary parts of $z$. Moreover, $\bar{z}=\Re(z)-\mathrm{i}\,\mathcal{I}(z)$ denotes the conjugate of $z$.
The space of bounded linear operators from $H$ to $H$ is denoted by $\mathcal{L}(H)$. The space of compact operators from $H$ to $H$ is denoted by $(K(H),\left\|\cdot\right\|_{\mathcal{L}(H)})$. The adjoint of an operator $A\in\mathcal{L}(H)$ is denoted by $A^{*}$, i.e., $\langle Ah,g\rangle_{H}=\langle h,A^{*}g\rangle_{H}$ for all $h\in H$. For $A\in\mathcal{L}(H^{\mathbb{C}})$ we define the transpose $A^{T}\in\mathcal{L}(H^{\mathbb{C}})$ by $\langle A^{T}h,g\rangle_{H^{\mathbb{C}}}=\langle A\bar{g},\bar{h}\rangle_{H^{\mathbb{C}}}$, $h,g\in H^{\mathbb{C}}$. Note that in contrast to $A^{*}$, the operator $A^{T}$ corresponds to the transpose without conjugation. Note that any $A\in\mathcal{L}(H)$ extends in a trivial and norm-conserving way to an operator $\tilde{A}\in\mathcal{L}(H^{\mathbb{C}})$ by setting $\tilde{A}(h+ig)=Ah+iAg$, $h,g\in H$.
For all $p\in[1,\infty)$ let $(\mathcal{L}_{p}(H),\left\|\cdot\right\|_{\mathcal{L}_{p}(H)})$ be the Banach space of Schatten class operators from $H$ to $H$, i.e.,
@@
and $\|A\|_{\mathcal{L}_{p}(H)}^{p}=\sum_{\lambda\in\sigma(A^{*}A)}\lambda^{p/2}$. In particular, we let $\mathcal{L}_{1}(H)$ and $\mathcal{L}_{2}(H)$ denote respectively the space of trace class operators and the space of Hilbert-Schmidt operators on $H$. Recall that $\mathcal{L}_{2}(H)$ is a Hilbert space when endowed with the inner product
@@
Writing $V^{\prime}$ for the dual of a Banach space $V$, we recall (see, e.g., [13, Section 19]) that the dual space of trace class operators satisfies $(\mathcal{L}_{1}(H_{1}))^{\prime}\simeq\mathcal{L}(H)$ under the duality paring

### v2

@@ Abstract.
We study the pricing of European-style options written on forward contracts within function-valued infinite-dimensional affine stochastic volatility models. The dynamics of the underlying forward price curves are modeled within the Heath-Jarrow-Morton-Musiela framework as solution to a stochastic partial differential equation modulated by a stochastic volatility process. We analyze two classes of affine stochastic volatility models: (i) a Gaussian model governed by a finite-rank Wishart process, and (ii) a pure-jump affine model extending the Barndorff–Nielsen–Shephard framework with state-dependent jumps in the covariance component. For the pure-jump model, we derive conditions for the existence of exponential moments and develop semi-closed Fourier-based pricing formulas for vanilla call and put options written on forward contracts with fixed delivery lag. For the Wishart model, we develop finite-rank Riccati approximations and corresponding Fourier-based pricing approximations. Our approach allows for tractable pricing in function-valued forward-curve models while capturing maturity-specific and term structure risk essential in forward markets.
@@ 1. Introduction
The aim of this work is to derive semi-closed pricing formulas for European vanilla options written on forward contracts in a function-valued HJMM model with stochastic volatility, and to validate these formulas numerically. Our pricing results are formulated for option families with a fixed delivery lag. To achieve this, we adopt the Heath–Jarrow–Morton–Musiela (HJMM) paradigm, originally proposed for interest-rate markets by [27] and later transferred to commodity markets by [11, 3]. In this paradigm, forward prices are modeled directly in time-to-maturity coordinates. Under this representation, the forward curve evolves according to a hyperbolic stochastic partial differential equation (SPDE) driven by a cylindrical Brownian motion whose covariance is itself modeled by a stochastic process that we call the instantaneous covariance process. We investigate two affine specifications of this covariance process: (i) infinite-dimensional Wishart processes taking values in the cone of positive self-adjoint trace class operators on a separable real Hilbert space, as introduced in [13], and (ii) a state-dependent jump extension of the operator-valued Ornstein–Uhlenbeck process introduced in [15], in which the jump intensity measure depends affinely on the current covariance state. The pure-jump model admits semi-explicit Fourier-Laplace expressions in terms of generalized Riccati equations, while for the Wishart model we work with finite-rank Riccati approximations. Exploiting this structure, together with establishing rigorous conditions for the existence of exponential moments in the pure-jump setting, enables the derivation of semi-closed pricing formulas for vanilla call and put options and finite-rank pricing approximations in the Wishart setting.
@@ 1.1. The HJMM framework
For any time $t\geq 0$ and maturity date $T\geq t$, the forward price $F(t,T)$ represents the agreed-upon price at time $t$ for a transaction occurring at the future date $T$. Utilizing the Musiela parametrization $x\equiv T-t$, forward prices become functions of time-to-maturity. We define the logarithmic forward prices as $f_{t}(x)\equiv\ln(F(t,t+x))$, where $x,t\geq 0$, and call the mapping $x\mapsto f_{t}(x)$ the (logarithmic) forward price curve at time $t\geq 0$. The forward curve dynamics are described by the stochastic partial differential equation (SPDE):
$\begin{cases}\mathop{}\!\mathrm{d}f_{t}(x)&=\left(\frac{\partial}{\partial x}f_{t}(x)+g_{t}(x)\right)\mathop{}\!\mathrm{d}t+\sum_{i=1}^{d}\sigma_{t}^{(i)}(x)\mathop{}\!\mathrm{d}W_{t}^{(i)},\\ f_{0}(x)&=\ln(F(0,x)),\end{cases}$ (1.1)
@@
The functions $(\sigma_{t}^{(i)}(x))_{0\leq t\leq T}$, ${i=1},\ldots,{d}$, are referred to as the instantaneous volatilities of the forward curve. They characterize the local covariance structure between forward prices of different maturities. The integer $d\in\mathbb{N}\cup\{+\infty\}$ is called the rank of the noise and corresponds to the number of independent risk factors driving the forward curves. If $d<\infty$, we refer to (1.1) as a finite-rank HJM model; otherwise, it is an infinite-rank HJM model.
@@ 1.2. Stochastic covariance models for forward curve dynamics
The diffusion term in (1.1) can be interpreted in a more abstract form as the stochastic integral of an operator-valued process $(\sigma_{t})_{0\leq t\leq T}$ acting on the increments of a cylindrical Brownian motion $(\mathop{}\!\mathrm{d}W_{t})_{0\leq t\leq T}$. We understand $(\sigma_{t})_{0\leq t\leq T}$ as the stochastic volatility operator, encoding the time-varying covariance structure of the forward curve dynamics. The functions $\sigma_{t}^{(i)}$ can therefore be interpreted as the principal components of $\sigma_{t}$, capturing the volatility structure along different directions in the underlying space of forward curves. The joint evolution of the forward curve and its stochastic volatility, represented by the process $(f_{t},\sigma_{t})_{0\leq t\leq T}$, thus defines an infinite-dimensional stochastic volatility model for forward curve dynamics.
The motivation for adopting a stochastic volatility framework stems from empirical evidence in commodity markets: volatility is time-varying, exhibits clustering, and is prone to abrupt spikes; see, e.g., [22, 5]. Classical models with constant volatility cannot capture these phenomena, whereas stochastic volatility models allow for a more realistic and flexible description of market behavior. Furthermore, to accurately capture volatility smiles in commodity option markets, incorporating stochastic volatility is essential. In contrast to approaches that perform a priori dimension reduction of the covariance, the pure-jump specification is formulated directly on the full maturity space rather than on a PCA-truncated factor model. For the Wishart specification considered here, the covariance state remains finite rank for fixed initial rank, and our analytical and numerical treatment is therefore based on finite-rank Riccati approximations. This still preserves the infinite-dimensional maturity state space while keeping the Gaussian covariance dynamics implementable. Empirical studies support the relevance of high-dimensional maturity dynamics: for instance, [1] and [32] demonstrate that power forward markets exhibit substantial idiosyncratic risk, and that high-dimensional noise is needed to capture the observed variability. In particular, the latter shows that even in PCA, around 10 factors are typically required to explain the majority of the variance, underlining the importance of accommodating rich maturity dynamics from the outset. Here our function-valued formulation helps, as it is essentially non-parametric, relying on the full initial forward curve and the full covariance operator as model inputs. The covariance structure is encoded via a positive kernel operator, which can capture empirical features with a minimal set of hyperparameters and enables direct calibration to market data, including option prices and historical term structure dynamics.
@@ 1.3. Contribution
In this paper, we contribute to this growing literature by developing a pricing framework for European options written on forward contracts under function-valued affine stochastic volatility. We formulate the forward curve dynamics within the HJMM framework and model the instantaneous covariance as an operator-valued affine process. The resulting risk-neutral pricing problem is treated for a fixed delivery lag. Two concrete model classes are considered: a Gaussian specification driven by a Wishart covariance process on the cone of positive trace-class operators, and a pure-jump model based on an extension of an operator-valued Ornstein–Uhlenbeck process to accommodate state-dependent jumps. The inclusion of jumps in the volatility process enhances the model’s ability to reflect sudden market shocks, which commonly arise from supply disruptions, geopolitical tensions, or abrupt changes in demand; see, e.g. [33]. For the pure-jump specification, we rigorously extend the affine transform formula to a complex strip in an infinite-dimensional Hilbert space and establish sufficient conditions under which exponential moments exist in real and complex domains. In particular, we identify a sufficient admissible domain for the moment generating function using generalized Riccati equations in Banach spaces, building on the analytic machinery introduced in [34], and extending prior finite-dimensional results such as those in [30]. For the Wishart specification, we work with finite-rank Riccati systems and the resulting Fourier-based pricing approximations.
Based on these results, we derive semi-closed-form pricing formulas for vanilla call and put options in the pure-jump model, together with finite-rank pricing approximations in the Wishart model, which enable numerical evaluation via Fourier inversion. We validate the accuracy and robustness of this pricing methodology through comparisons with direct or conditional Monte Carlo benchmarks. The numerical results show that the first forward moment in the Wishart experiment, the corrected Wishart option prices, and the option prices in the pure-jump experiments are consistent with these benchmarks, see Figures 1, 2, 3, and 4 below. For the explicit Lévy and BNS specifications, the affine formulas provide substantial computational gains over Monte Carlo simulation. In the state-dependent jump example, the relative runtime ranking is much more implementation-dependent and should therefore not be overinterpreted. Our results demonstrate that, despite the infinite-dimensional nature of the maturity state space, affine stochastic volatility models remain useful tools for capturing rich dynamics in forward and futures markets, provided the tractability claims are interpreted in line with the particular specification under consideration.
@@ 1.4. Layout of the article
The remainder of the paper is organized as follows. In Section 2, we introduce the modeling framework for forward curve dynamics using the HJMM approach and define the class of infinite-dimensional stochastic volatility models under consideration. Sections 2.1 and 2.2 present two concrete affine specifications for the volatility process: a pure-jump model based on an extension of an operator-valued Ornstein–Uhlenbeck process, and a Gaussian model driven by a Wishart process on the cone of positive trace-class operators. In Section 3, we provide Fourier-based pricing formulas for the pure-jump model and finite-rank pricing approximations for the Wishart model. Section 4 presents a numerical study comparing these affine formulas with Monte Carlo simulation methods. We highlight the accuracy, robustness, and computational efficiency of the affine approach, especially in settings where simulation becomes numerically challenging. Finally, in Section 5, we study the existence of the exponential moments.
@@ 1.5. Notation
For $(X,\tau)$ a topological space and $S\subset X$ we let $\mathcal{B}(S)$ denote the Borel-$\sigma$-algebra generated by the relative topology on $S$. We denote by $C^{k}([0,T];S)$ the space of $S$-valued $k$-times continuously differentiable functions on $[0,T]$. Throughout this article we fix a separable, infinite-dimensional real Hilbert space $(H,\langle\cdot,\cdot\rangle_{H})$. The complexification of a real Hilbert space is the complex vector space $H^{\mathbb{C}}=H\oplus_{\mathbb{R}}\mathrm{i}H=\{x+\mathrm{i}y\colon x,y\in H\}$ equipped with the complex bilinear extension of the real inner product, $\langle x+\mathrm{i}y,u+\mathrm{i}v\rangle_{H^{\mathbb{C}}}=\langle x,u\rangle_{H}-\langle y,v\rangle_{H}+\mathrm{i}\langle y,u\rangle_{H}+\mathrm{i}\langle x,v\rangle_{H}$, for all $x,y,u,v\in H$. Note that $\langle\cdot,\cdot\rangle_{H^{\mathbb{C}}}$ is bilinear (not sesquilinear) and hence not a Hilbert-space inner product; it is the natural pairing for the analytic continuation of the Fourier–Laplace transform used throughout this paper. For $z=x+\mathrm{i}y\in H^{\mathbb{C}}$, we say that $x=\Re(z)$ and $y=\mathcal{I}(z)$ are the real and imaginary parts of $z$. Moreover, $\bar{z}=\Re(z)-\mathrm{i}\,\mathcal{I}(z)$ denotes the conjugate of $z$.
The space of bounded linear operators from $H$ to $H$ is denoted by $\mathcal{L}(H)$. The space of compact operators from $H$ to $H$ is denoted by $(K(H),\left\|\cdot\right\|_{\mathcal{L}(H)})$. The adjoint of an operator $A\in\mathcal{L}(H)$ is denoted by $A^{*}$, i.e., $\langle Ah,g\rangle_{H}=\langle h,A^{*}g\rangle_{H}$ for all $h\in H$. For $A\in\mathcal{L}(H^{\mathbb{C}})$ we define the transpose $A^{T}\in\mathcal{L}(H^{\mathbb{C}})$ by $\langle A^{T}h,g\rangle_{H^{\mathbb{C}}}=\langle A\bar{g},\bar{h}\rangle_{H^{\mathbb{C}}}$, $h,g\in H^{\mathbb{C}}$. Note that in contrast to $A^{*}$, the operator $A^{T}$ corresponds to the transpose without conjugation. Note that any $A\in\mathcal{L}(H)$ extends in a trivial and norm-conserving way to an operator $\tilde{A}\in\mathcal{L}(H^{\mathbb{C}})$ by setting $\tilde{A}(h+ig)=Ah+iAg$, $h,g\in H$. For all $p\in[1,\infty)$ let $(\mathcal{L}_{p}(H),\left\|\cdot\right\|_{\mathcal{L}_{p}(H)})$ be the Banach space of Schatten class operators from $H$ to $H$, i.e.,
@@
and $\|A\|_{\mathcal{L}_{p}(H)}^{p}=\sum_{\lambda\in\sigma(A^{*}A)}\lambda^{p/2}$, where eigenvalues are counted with multiplicity. In particular, we let $\mathcal{L}_{1}(H)$ and $\mathcal{L}_{2}(H)$ denote respectively the space of trace class operators and the space of Hilbert-Schmidt operators on $H$. Recall that $\mathcal{L}_{2}(H)$ is a Hilbert space when endowed with the inner product
@@
Writing $V^{\prime}$ for the dual of a Banach space $V$, we recall (see, e.g., [12, Section 19]) that the dual space of trace class operators satisfies $(\mathcal{L}_{1}(H))^{\prime}\simeq\mathcal{L}(H)$ under the duality pairing